methylphenidate (ritalin)is recommended 0.3-2 mg/kg/day. What is the maximum recommended dose per day in mg a child who weighs 66lb?
Answer #1

ANSWER.

* Calculation of mg of drug required for a child weight 66lb.

          Recommended dose = 0.3 - 2 mg/kg/day.

          Weight of the patient = 66lb.

     Convert lb to kg for calculation

                   1lb = 0.454 kg

        Weight of the patient in kg = 66 * 0.454

                                                   = 29.9 kg.

    So

         mg of drug required for this child per day = 0.3 * 29.9 - 2 * 29.9

                                                                           = 8.9 mg to 59.8 mg per day